Translate the following statement into an algebraic equation. Let x represent the number. 8less than 2times a numberis 9 times that same number.
Found 2 solutions by funmath, checkley71:
Answer by funmath(2873)

Translate the following statement into an algebraic equation. Let x represent the number. 8 less than 2 times a number is 9 times that same number.
"8 less" means -8
"2 times" means 2*
"is" means =
"9 times" means 9*
The expression is:
2*x-8=9*x
2x-8=9x
Happy Calculating!!!

Answer by checkley71(8403)

2X-8=9X
2X-9X=8
-7X=8
X=-8/7
PROOF
2*-8/7-8=9*-8/7
-16/7-8=-72/7
-16/7-56/7=-72/7
-72/7=-72/7